# Tokenizing Buildings: A Transformer for Layout Synthesis

### v1

@@ Abstract
We introduce Small Building Model (SBM), a Transformer-based architecture for layout synthesis in Building Information Modeling (BIM) scenes. We address the question of how to tokenize buildings by unifying heterogeneous feature sets of architectural elements into sequences while preserving compositional structure. Such feature sets are represented as a sparse attribute-feature matrix that captures room properties. We then design a unified embedding module that learns joint representations of categorical and possibly correlated continuous feature groups. Lastly, we train a single Transformer backbone in two modes: an encoder-only pathway that yields high-fidelity room embeddings, and an encoder–decoder pipeline for autoregressive prediction of room entities—referred to as Data-Driven Entity Prediction (DDEP). Experiments across retrieval and generative layout synthesis show that SBM learns compact room embeddings that reliably cluster by type and topology, enabling strong semantic retrieval. In DDEP mode, SBM produces functionally sound layouts—with fewer collisions and boundary violations and improved navigability.
@@ 1 Introduction
Professional Computer Aided Design (CAD)/BIM scenes encode rich semantics, hierarchies, and domain constraints, but most strong 3D generators—voxel, mesh, point-cloud, or image-conditioned diffusion models—treat scenes as unstructured geometry [eastman_spatial_1975], yielding visually plausible yet hard-to-edit outputs that frequently violate basic validity rules. Room-scale layout design further demands multi-scale reasoning and parametric outputs compatible with downstream BIM tools, reflecting the highly constrained yet repetitive nature of practice, where designers position walls, doors, casework, and circulation to satisfy codes and programs across hundreds of similar rooms [alphabet_carpo, carpo2017second, eastman_use_1975, Flager2007ACO]. Decades of computational approaches—from CAD macros and parametric families to rule systems, optimization, and learning-based methods—have aimed to automate this process [MONEDERO2000369], but progress remains bottlenecked not by algorithms themselves but by the representations they operate on. An effective representation must expose structure, generalize across building types, and maintain BIM-level geometric and semantic coherence, underscoring the need to rethink spatial data encoding as the foundation for scalable layout reasoning.
To this end, we introduce a normalized hierarchical tokenizer that encodes room topology, entity attributes, wall-referenced geometry, and relational structure into compact BIM-Token Bundles. We realize these bundles as rows of a sparse attribute–feature matrix and embedded via a mixed-type module that produces a unified token vector for Transformer models. Building on this representation, our SBM provides a Transformer backbone for BIM scenes, supporting both encoder-only analysis of room embeddings at scale and encoder–decoder generation through DDEP, which autoregressively places room-level entities with mixed categorical and continuous attributes while preserving the compositional structure of BIM data.
In summary, we contribute (1) a normalized hierarchical tokenization of BIM scenes with a mixed-type embedding module, (2) a unified Transformer backbone (SBM) that supports both retrieval and generative layout synthesis through encoder-only and encoder–decoder (DDEP) modes, and (3) a comprehensive evaluation on a professional BIM dataset demonstrating consistently higher scene coherence, geometric validity, and constraint satisfaction than recent LLM/VLM and domain-specific baselines.
@@ 2 Related Work
Other systems use LLMs for guided refinement: HouseTune [zong2024housetune] proposes and linguistically adjusts floorplan candidates, while FlairGPT [littlefair2025flairgpt] aligns language tokens with furniture and style attributes for interior design exploration. Multi-modal models like LLM4CAD [li2024llm4cad] tokenize CAD elements symbolically to support reasoning but operate at the object level without encoding continuous spatial attributes or wall-referenced coordinates. Recently, FloorPlan-DeepSeek [yin2025floorplandeepseek] formulates floorplan synthesis as autoregressive next-room prediction, encoding each room as a vector of semantic type and geometric attributes. This sequential approach constructs layouts incrementally from partial plans or text prompts, achieving improved structural coherence and enabling downstream editing. However, the approach remains limited to room-level floor-plan arrangement, and does not encode fine-grained BIM elements or wall-hosted entities.
Gaps.
Spatial layout research has evolved from rule-based systems to graphs, images, vectorized polygons, and most recently language-based token sequences, each step improving expressiveness. Yet existing representations remain rigid, task-specific, or fragmented, and still fail to unify semantic, relational, and geometric information in a sequence-native form. Most methods yield images, semantic maps, or coarse graphs rather than structured tokens with explicit geometric references and mixed categorical–continuous attributes, limiting constraint-aware reasoning, cross-task generalization, and downstream BIM editability.
This persistent representational gap motivates our approach: a normalized tokenization of BIM scenes coupled with a custom embedding module and a shared Transformer backbone, designed to provide a unified, expressive, and geometrically grounded foundation for scalable spatial layout reasoning.
@@ 3.1 Problem Setup and BIM Room Representation
We operate on room-level layouts extracted from professional BIM projects (see Sec. 4.1.1 for dataset details). Each room is treated as a parametric layout with semantic attributes, and wall, openings, and furniture/casework entities, all expressed in a local coordinate frame normalized for position and scale.
@@
We explicitly decompose each room into a room envelope—the walls, openings, and coarse layout attributes—and a set of room contents (props and casework). We write $r_{\text{env}}=\big(y^{\text{topo}},y^{\text{layout}},\mathcal{E},\mathcal{D},\mathcal{W}\big),r_{\text{ent}}=\big(\mathcal{P},\mathcal{C}\big)$ and define the full room as $r=(r_{\text{env}},r_{\text{ent}})$. Here $y^{\text{topo}}$ is a discrete room-type token (e.g., bedroom, kitchen) and $y^{\text{layout}}$ is a layout token that aggregates global scalar attributes such as area and perimeter. The sets $\mathcal{E},\mathcal{D},\mathcal{W}$ are walls, doors, and windows, and $\mathcal{P},\mathcal{C}$ are props and casework, respectively. In our model, the encoder operates only on the envelope $r_{\text{env}}$, while the decoder operates only on the entity sequence $r_{\text{ent}}$ conditioned on the encoder output.
@@
Doors and windows are modeled as objects attached to specific wall edges with wall-referenced coordinates. For doors we write $\mathcal{D}=\{d_{k}\}_{k=1}^{N_{D}},d_{k}=\big(j_{k},t_{k},w_{k},a^{\text{door}}_{k}\big),$ where $j_{k}\in\{1,\dots,N_{E}\}$ indexes the supporting edge $e_{j_{k}}$, $t_{k}\in[0,1]$ is a normalized position along that edge, $w_{k}$ is the opening width, and $a^{\text{door}}_{k}$ includes additional categorical attributes (e.g., door family/type, swing direction). Windows $\mathcal{W}=\{w_{\ell}\}_{\ell=1}^{N_{W}}$ are defined analogously with edge attachment, normalized position, width, and window-specific attributes. This wall-referenced parameterization makes openings invariant to absolute translation and scale and directly compatible with BIM editing operations.
@@
Room contents are represented as two sets of entities: $\mathcal{P}=\{p_{u}\}_{u=1}^{N_{P}},\qquad\mathcal{C}=\{c_{v}\}_{v=1}^{N_{C}},$ for props (furniture) and casework. Each entity is parameterized by a discrete type and a small set of wall-referenced continuous attributes. For a generic entity $q\in\mathcal{P}\cup\mathcal{C}$ we write $q=\big(c_{q},j_{q},t_{q},\delta_{q},s_{q},\rho_{q},a^{\text{ent}}_{q}\big),$ where $c_{q}$ is an entity type identifier (e.g., bed, base_cabinet), $j_{q}$ is the supporting edge index, $t_{q}\in[0,1]$ is the position along that edge, $\delta_{q}$ is a signed lateral offset from the wall into the room, $s_{q}$ encodes size or footprint parameters (e.g., width/depth), $\rho_{q}$ is an in-plane rotation angle when applicable (props only), and $a^{\text{ent}}_{q}$ contains any additional categorical attributes. This representation captures both semantic identity and precise placement of entities relative to the room envelope.
@@
Given this decomposition, our model supports two tasks. In encoder-only mode, the goal is to map the room envelope $r_{\text{env}}$ to a compact embedding $z(r_{\text{env}})\in\mathbb{R}^{d}$ that preserves semantic and geometric relationships between rooms (for retrieval and analysis). In encoder–decoder mode (DDEP), the encoder consumes only the envelope $r_{\text{env}}$ to produce a contextual memory, and the decoder autoregressively generates the entity sequence $r_{\text{ent}}=(\mathcal{P},\mathcal{C})$ as a sequence of tokens (see Fig. 2). The following subsections describe how we convert $r_{\text{env}}$ and $r_{\text{ent}}$ into BIM-Token Bundles on sparse attribute–feature matrices, embed these mixed-type features, and train a shared Transformer backbone to solve both tasks.
@@ 3.2 BIM Tokenization
Given the room envelope $r_{\text{env}}$ and contents $r_{\text{ent}}$ (Sec. 3.1), we convert each room into two ordered sequences of BIM-Token Bundles: an envelope sequence for the encoder and an entity sequence for the decoder. Each bundle corresponds to a single logical element (topology, layout, wall, door, window, or entity) and aggregates all of its semantic and geometric attributes into a sparse feature vector.
@@
On the decoder side, we form a contents-only entity sequence $\big(\tau^{\text{SOS}},\tau^{\text{ent}}_{1},\dots,\tau^{\text{ent}}_{T},\tau^{\text{EOS}}\big),$ where each $\tau^{\text{ent}}_{t}$ corresponds to one entity $q\in\mathcal{P}\cup\mathcal{C}$ with its type and wall-referenced attributes, and $\tau^{\text{SOS}}$, $\tau^{\text{EOS}}$ are special start/end tokens. The room envelope is provided to the decoder only via cross-attention to the encoder memory, not re-encoded as decoder tokens.
@@
We construct the decoder matrix $X^{\text{dec}}\in\mathbb{R}^{F_{\text{dec}}\times S_{\text{dec}}}$ in the same way, now with decoder-specific features such as entity type, edge attachment, wall-referenced coordinates $(t_{q},\delta_{q})$, size parameters $s_{q}$, and rotation $\rho_{q}$, plus token_type_id and token_id. Here too, only the rows relevant to each token type are active and the rest are set to p=100p=-100. In the next subsection, we describe how these sparse matrices $X^{\text{enc}}$ and $X^{\text{dec}}$ are mapped to dense token embeddings by a mixed-type embedding module.
@@ 3.3 Mixed-Type Embedding Module
Let $\mathcal{F}_{\text{enc}}$ be the set of encoder features. For each feature $f\in\mathcal{F}_{\text{enc}}$ and token position $s$, we define a valid-feature mask $m_{f,s}=\mathds{1}\big[X^{\text{enc}}_{f,s}\neq-100\,],$ which indicates whether feature $f$ applies to token $s$. After replacing sentinel values 100-100 by 0 for numerical stability inside $E_{f}$, we compute feature embeddings and apply the mask: $u_{f,s}=E_{f}\big(X^{\text{enc}}_{f,s}\big)\in\mathbb{R}^{d},\tilde{u}_{f,s}=m_{f,s}\cdot u_{f,s}.$ This second masking step ensures that inactive features contribute exactly zero, including the bias terms of linear layers. The encoder attention mask further zeroes out all embeddings for padded token positions.
@@
By treating each BIM-Token Bundle as a sparse feature column and mapping it through this mixed-type embedding module, we obtain dense, sequence-native token embeddings for both the room envelope and its contents. The next subsection describes how a standard Transformer backbone consumes $E^{\text{enc}}$ and $E^{\text{dec}}$ in encoder-only and encoder–decoder modes.
@@ 3.4 Transformer Backbone and Operating Modes
Given the encoder and decoder token embeddings we employ a standard Transformer encoder–decoder architecture. The same backbone is used in two operating modes: encoder-only (for room embeddings and retrieval) and encoder–decoder (for data-driven entity prediction, DDEP).
@@
In DDEP mode, we use the full encoder–decoder Transformer to generate the room contents $r_{\text{ent}}=(\mathcal{P},\mathcal{C})$ conditioned on the room envelope $r_{\text{env}}$. The encoder processes the envelope as above, producing a memory $M$ that is held fixed during decoding. The decoder takes entity token embeddings $E^{\text{dec}}$ and attends both to the prefix of generated tokens (causal self-attention) and to the encoder memory $M$ (cross-attention), $H=\mathrm{Dec}_{\theta}\big(E^{\text{dec}},M\big)\in\mathbb{R}^{S_{\text{dec}}\times d},$ where $\mathrm{Dec}_{\theta}$ is a stack of masked self-attention, cross-attention, and feed-forward layers. Decoder states in $H$ feed into multi-head prediction layers that produce mixed categorical and continuous outputs for each entity token. We enforce a constrained structured decoding schema which refines the raw predictions.

### v2

@@ Abstract
We introduce Small Building Model (SBM), a Transformer-based architecture for layout synthesis in Building Information Modeling (BIM) scenes. We address the question of how to tokenize buildings by unifying heterogeneous feature sets of architectural elements into sequences while preserving compositional structure. Such feature sets are represented as a sparse attribute-feature matrix that captures room properties. We then design a unified embedding module that learns joint representations of categorical and possibly correlated continuous feature groups. Lastly, we train a single Transformer backbone in two modes: an encoder-only pathway that yields high-fidelity room embeddings, and an encoder–decoder pipeline for autoregressive prediction of residential room entities—referred to as Data-Driven Entity Prediction (DDEP). Experiments across retrieval and generative layout synthesis show that SBM learns compact room embeddings that reliably cluster by type and topology, enabling strong semantic retrieval. In DDEP mode, SBM produces functionally sound layouts—with fewer collisions and boundary violations, and improved navigability—outperforming general-purpose LLM/VLM baselines and recent domain-specific methods. A controlled same-data benchmark confirms that these gains are architectural, not solely data-driven. An anonymized version of our dataset and code will be released upon acceptance.
@@ 1 Introduction
Professional Computer Aided Design (CAD)/BIM scenes encode rich semantics, hierarchies, and domain constraints, but most strong 3D generators—voxel, mesh, point-cloud, or image-conditioned diffusion models—treat scenes as unstructured geometry [eastman_spatial_1975], yielding visually plausible yet hard-to-edit outputs that frequently violate basic validity rules. Room-scale layout design further demands multi-scale reasoning and parametric outputs compatible with downstream BIM tools, reflecting the highly constrained yet repetitive nature of practice, where designers position walls, doors, casework, and circulation to satisfy codes and programs across hundreds of similar rooms [alphabet_carpo, carpo2017second, eastman_use_1975, Flager2007ACO]. Decades of computational approaches—from CAD macros and parametric families to rule systems, optimization, and learning-based methods—have aimed to automate this process [MONEDERO2000369], but progress remains bottlenecked not by algorithms themselves but by the representations they operate on. An effective representation must expose structure, generalize across typologies, and maintain BIM-level geometric and semantic coherence, underscoring the need to rethink spatial data encoding as the foundation for scalable layout reasoning.
To this end, we introduce a normalized hierarchical tokenizer that encodes room topology, entity attributes, wall-referenced geometry, and relational structure into compact BIM-Token Bundles. We realize these bundles as rows of a sparse attribute–feature matrix and embedded via a mixed-type module that produces a unified token vector for Transformer models. Building on this representation, our SBM provides a Transformer backbone for BIM scenes, supporting both encoder-only understanding of room embeddings at scale and encoder–decoder generation through DDEP, which autoregressively places room-level entities with mixed categorical and continuous attributes while preserving the compositional structure of BIM data.
In summary, we contribute (1) a normalized hierarchical tokenization of residential BIM scenes with a mixed-type embedding module, (2) a unified Transformer backbone (SBM) that supports both retrieval and generative layout synthesis through encoder-only and encoder–decoder (DDEP) modes, and (3) a two-track evaluation—a frontier LLM benchmark and a controlled same-data benchmark with ablations—demonstrating that DDEP’s advantages in scene coherence, geometric validity, and constraint satisfaction are architectural, not solely data-driven.
@@ 2 Related Work
Transformer-based indoor-scene synthesis (object-centric).
A closely related line of work treats indoor scene synthesis as Transformer-driven generation of object instances (category and pose) conditioned on a room boundary or floor-plan mask. ATISS [Paschalidou2021NEURIPS] and SceneFormer [wang2020sceneformer] generate sequences (or sets) of furniture objects with discretized/normalized geometric attributes, targeting visually plausible and collision-free arrangements in synthetic scene datasets. CLIP-Layout [Liu2023CLIPLayoutSI] further augments this object-centric paradigm with semantic furniture embeddings (via CLIP) to encourage style consistency and improve instance-level retrieval during generation. While highly relevant, these approaches largely assume a flat object vocabulary (furniture instances with poses) and do not model the heterogeneous, hierarchical, and parametric nature of BIM: envelopes and walls, wall-hosted elements, typed constraints, and mixed categorical–continuous attributes that must remain editable in downstream CAD/BIM workflows. This mismatch motivates BIM-native tokenization and embedding that preserve compositional structure (e.g., envelope $\rightarrow$ boundaries $\rightarrow$ hosted entities) and enable constraint-aware reasoning beyond object boxes.
In parallel, controllable 2D layout generation for graphic design has been studied with Transformer formulations such as BLT [10.1007/978-3-031-19790-1_29], which uses bidirectional/masked prediction and iterative refinement to satisfy user constraints while accelerating inference.
@@
Other systems use LLMs for guided refinement: HouseTune [zong2024housetune] proposes and linguistically adjusts floorplan candidates, while FlairGPT [littlefair2025flairgpt] aligns language tokens with furniture and style attributes for interior design exploration. Multi-modal models like LLM4CAD [li2024llm4cad] tokenize CAD elements symbolically to support reasoning but operate at the object level without encoding continuous spatial attributes or wall-referenced coordinates. Recently, FloorPlan-DeepSeek [yin2025floorplandeepseek] formulates floorplan synthesis as autoregressive next-room prediction, encoding each room as a vector of semantic type and geometric attributes. This sequential approach constructs layouts incrementally from partial plans or text prompts, achieving improved structural coherence and enabling downstream editing. However, the approach remains limited to room-level floor-plan arrangement, and does not encode fine-grained BIM elements or wall-hosted entities. Across this progression, existing representations still fail to unify semantic, relational, and geometric information in a sequence-native form with mixed categorical–continuous attributes, limiting constraint-aware reasoning and downstream BIM editability. This gap motivates our normalized BIM tokenization coupled with a shared Transformer backbone.
@@ 3.1 Problem Setup and BIM Room Representation
We operate on residential room-level layouts extracted from professional BIM projects (see Sec. 4 for dataset details), expressed in a local coordinate frame normalized for position and scale.
@@
Each room is decomposed into a room envelope—walls, openings, and coarse layout attributes—and room contents (props and casework): $r_{\text{env}}=\big(y^{\text{topo}},y^{\text{layout}},\mathcal{E},\mathcal{D},\mathcal{W}\big),r_{\text{ent}}=\big(\mathcal{P},\mathcal{C}\big)$ and define the full room as $r=(r_{\text{env}},r_{\text{ent}})$. Here $y^{\text{topo}}$ is a discrete room-type token (e.g., bedroom, kitchen) and $y^{\text{layout}}$ is a layout token that aggregates global scalar attributes such as area and perimeter. The sets $\mathcal{E},\mathcal{D},\mathcal{W}$ are walls, doors, and windows, and $\mathcal{P},\mathcal{C}$ are props and casework, respectively. In our model, the encoder operates only on the envelope $r_{\text{env}}$, while the decoder operates only on the entity sequence $r_{\text{ent}}$ conditioned on the encoder output.
@@
Doors and windows are modeled as objects attached to specific wall edges with wall-referenced coordinates. For doors we write $\mathcal{D}=\{d_{k}\}_{k=1}^{N_{D}},d_{k}=\big(j_{k},t_{k},w_{k},a^{\text{door}}_{k}\big),$ where $j_{k}\in\{1,\dots,N_{E}\}$ indexes the supporting edge $e_{j_{k}}$, $t_{k}\in[0,1]$ is a normalized position along that edge, $w_{k}$ is the opening width, and $a^{\text{door}}_{k}$ includes additional categorical attributes (e.g., door family/type, swing direction). Windows $\mathcal{W}=\{w_{\ell}\}_{\ell=1}^{N_{W}}$ are defined analogously. This wall-referenced parameterization makes openings invariant to absolute translation and scale.
@@
Room contents are represented as two sets of entities: $\mathcal{P}=\{p_{u}\}_{u=1}^{N_{P}},\qquad\mathcal{C}=\{c_{v}\}_{v=1}^{N_{C}},$ for props (furniture) and casework. Each entity is parameterized by a discrete type and a small set of wall-referenced continuous attributes. For a generic entity $q\in\mathcal{P}\cup\mathcal{C}$ we write $q=\big(c_{q},j_{q},t_{q},\delta_{q},s_{q},\rho_{q},a^{\text{ent}}_{q}\big),$ where $c_{q}$ is an entity type identifier (e.g., bed, base_cabinet), $j_{q}$ is the supporting edge index, $t_{q}\in[0,1]$ is the position along that edge, $\delta_{q}$ is a signed lateral offset from the wall, $s_{q}$ encodes size parameters (e.g., width/depth), $\rho_{q}$ is an in-plane rotation (props only), and $a^{\text{ent}}_{q}$ contains additional categorical attributes.
@@
Given this decomposition, our model supports two tasks. In encoder-only mode, the goal is to map the room envelope $r_{\text{env}}$ to a compact embedding $z(r_{\text{env}})\in\mathbb{R}^{d}$ that preserves semantic and geometric relationships between rooms (for retrieval and analysis). In encoder–decoder mode (DDEP), the encoder consumes only the envelope $r_{\text{env}}$ to produce a contextual memory, and the decoder autoregressively generates the entity sequence $r_{\text{ent}}=(\mathcal{P},\mathcal{C})$ as a sequence of tokens (see Fig.˜2).
@@ 3.2 BIM Tokenization
Given the room envelope $r_{\text{env}}$ and contents $r_{\text{ent}}$ (Sec.˜3.1), we convert each room into two ordered sequences of BIM-Token Bundles: an envelope sequence for the encoder and an entity sequence for the decoder. Each bundle corresponds to a single logical element (topology, layout, wall, door, window, or entity) and aggregates all of its semantic and geometric attributes into a sparse feature vector.
@@
On the decoder, we form a content only entity sequence $\big(\tau^{\text{SOS}},\tau^{\text{ent}}_{1},\dots,\tau^{\text{ent}}_{T},\tau^{\text{EOS}}\big),$ where each $\tau^{\text{ent}}_{t}$ corresponds to one entity $q\in\mathcal{P}\cup\mathcal{C}$ with its type and wall-referenced attributes, and $\tau^{\text{SOS}}$, $\tau^{\text{EOS}}$ are special start/end tokens. The room envelope is provided to the decoder only via cross-attention to the encoder memory, not re-encoded as decoder tokens.
@@
We construct the decoder matrix $X^{\text{dec}}\in\mathbb{R}^{F_{\text{dec}}\times S_{\text{dec}}}$ in the same way, now with decoder-specific features such as entity type, edge attachment, wall-referenced coordinates $(t_{q},\delta_{q})$, size parameters $s_{q}$, and rotation $\rho_{q}$, plus token type id and token id. Here too, only the rows relevant to each token type are active and the rest are set to p=100p=-100.
@@ 3.3 Mixed-Type Embedding Module
Let $\mathcal{F}_{\text{enc}}$ be the set of encoder features. For each feature $f\in\mathcal{F}_{\text{enc}}$ and token position $s$, we define a valid-feature mask $m_{f,s}=\mathds{1}\big[X^{\text{enc}}_{f,s}\neq-100\,],$ which indicates whether feature $f$ applies to token $s$. We then compute feature embeddings and apply the mask: $u_{f,s}=E_{f}\big(X^{\text{enc}}_{f,s}\big)\in\mathbb{R}^{d},\tilde{u}_{f,s}=m_{f,s}\cdot u_{f,s}.$ This second masking step ensures that inactive features contribute exactly zero, including the bias terms of linear layers.
@@ 3.4 Transformer Backbone and Operating Modes
Given the token embeddings we employ a standard Transformer encoder–decoder architecture. The same backbone is used in two operating modes: encoder-only (for room embeddings and retrieval) and encoder–decoder (for data-driven entity prediction, DDEP).
@@
In DDEP mode, we use the full encoder–decoder Transformer to generate the room contents $r_{\text{ent}}=(\mathcal{P},\mathcal{C})$ conditioned on the room envelope $r_{\text{env}}$. The encoder processes the envelope as above, producing a memory $M$ that is held fixed during decoding. The decoder takes entity token embeddings $E^{\text{dec}}$ and attends both to the prefix of generated tokens (causal self-attention) and to the encoder memory $M$ (cross-attention), $H=\mathrm{Dec}_{\theta}\big(E^{\text{dec}},M\big)\in\mathbb{R}^{S_{\text{dec}}\times d},$ where $\mathrm{Dec}_{\theta}$ is a stack of masked self-attention, cross-attention, and feed-forward layers. Decoder states in $H$ feed into multi-head prediction layers that produce mixed categorical and continuous outputs for each entity token. At inference time, a constrained decoding layer enforces BIM validity constraints (e.g., egress clearances, door-swing zones) by filtering the model’s output distribution at each step, ensuring structurally valid token sequences without external guidance (details in Suppl.).
Training.
We adopt a two-stage schedule. First, the encoder is pre-trained on the full corpus with a composite loss combining room-type classification, masked-token modeling, graded triplet contrastive learning, and geometric preservation. Second, the encoder–decoder model is fine-tuned for DDEP using teacher forcing: at each step the decoder predicts the next entity token given the ground-truth prefix, with a combined cross-entropy loss over categorical heads and MSE loss over continuous regression heads. Full loss formulations, hyperparameters, and infrastructure details are provided in the Supplementary Material.